3. Assume the required reserve ratio is 40% and the Open Market Committee of the FED sells $500 billion in bonds to the public. Assuming banks give out as many loans as possible, what is the total change in the money supply? What is the total change in Transaction Accounts? If the M1 was originally $9500 billion, what is the new M1 ( After the change)You must show your work.

4. Assume the required reserve ratio is 10% and the Open Market Committee of the FED buys $50 billion in bonds from the public. Assuming banks give out as many loans as possible, what is the total change in the money supply? What is the total change in Transaction Accounts? If the M1 was originally $9500 billion, what is the new M1 ( After the change)?You must show your work.
Answer #1

3):-An open market sale decreases money supply.

Decrease in money supply ($ billion) = Open = market sale / Reserve ratio = 500 / 0.4 = 1250

(ii) An open market sale decreases transaction accounts. Decrease in transaction accounts ($ billion) = = ease in money supply Open market sale = 1250 - 500 = 750

(iii) New M1 ($ billion) = 9500-1250 = 8250

According to HOMEWORKLIB RULES we have right to do one question in one time i hope you understand